# Space science & the space economy

### v1

@@ 5 Conclusions
(1) broaden the funding base. In addition to government money, scientists can fund their projects through philanthropic foundations, public-private partnerships, and private investors. Knowledge is a crucial space resource, and it is precisely the resource that basic science researchers are best at harvesting. Accurate Position, Navigation, and Timing, the largest business in today’s and also tomorrow’s space economy, moving hundreds billion dollars, is famously impossible without general relativity corrections. Paradoxically, Albert Einsten (the inventor of general relativity more than 100 years ago) and his heirs did not make a single dollar from this invention. More relevant to new space is the story of Riccardo Giacconi, 2002 Nobel Prize winner for the discovery of Cosmic X-ray sources and the Cosmic X-ray Background. When Giacconi joined Bruno Rossi of MIT in 1959 as a young researcher, he was hired by American Science and Engineering (AS&E), a company started the year before by Martin Annis and of which Bruno Rossi was chair of the board of directors. AS&E was working on NASA contracts, but was also very active in the market. They developed the first body scanner, later adopted in airports around the world. Money earned in the market allowed AS&E to let Rossi and Giacconi develop the strategies and instrumentation that led to the creation of X-ray astronomy and to a series of world-leading missions: Uhuru, Einstein Observatory, and Chandra [25]. This is one of the several possible scenarios that could link science and business. Today, the new space economy probably provides more opportunities than in the time of Rossi and Giacconi. Two broad areas seem particularly promising today and merit a mention: quantum technologies and resource prospecting on the Moon, Mars, and asteroids.
@@
(2) the ability of scientists to keep up with the innovations of the new space economy. Technology innovation proceeds through alternating phases of incremental and disruptive development. Business innovates through vertical integration and scale production, but also a service-oriented business model and a culture of innovation that manifest itself as openness to risk and iterative development. Giacconi again comes to our aid to emphasize these concepts. Soon after the discovery of a strong, seemingly uniform, X-ray background in 1962 Giacconi understood that to detect the sources responsible for the Cosmic X-ray background, an imaging quality of the order of an arcsecond was needed in X-rays, several thousands times better than then available. At those early times, it was difficult, if not impossible, to convince NASA to embark directly on a project to develop and launch an X-ray observatory with arcsecond capabilities. Instead, Giacconi imagined an iterative program aimed at consolidating the field and arriving at the arcsecond telescope (Chandra) through steps, Uhuru first, to provide the first all-sky survey of bright X-ray sources at the beginning of the 1970s, Einstein Observatory, to demonstrate in space the capabilities of X-ray optics, and finally the Advanced X-ray Astrophysics Facility, AXAF, renamed Chandra after launch in 1999. The new space economy offers the opportunity to perform iterative approach programs of this kind on much shorter timescale and at a much lower cost.
@@
We have searched the Astrophysical Data System (ADS) and the SCOPUS database for referred publications, citations, and the H index of a sample of recent space scientific missions funded by NASA and ESA. We included in the list a sample of cubesats realized by Universities and scientific institutes. The search was carried out using the simplest possible keywords, to be able to have a data sample as homogeneous as possible. As an example, the Gaia ADS Library produced by ESA includes 12842 refereed papers since 2014. The library is obtained by merging the results of a tailored and complex ADS query with a manually curated library. The authors claim a fraction of false positives $<$5%. The much simpler ADS query ((facility:Gaia OR abs:Gaia OR title:Gaia) year:2014-2024 property:refereed collection:astronomy) produces 6617 entries, of which $>$90% are present in the ESA Gaia ADS Library. A manual check on the papers missing in the ESA Gaia ADS Library shows that a large fraction are actually using Gaia data or results obtained from Gaia data. So we are confident that the fraction of false positives in our table is certainly smaller than 10% and probably smaller than 5%. Limiting the sample to papers which explicitly mention the name of the mission in the title, abstract, or facility keyword would include most papers that are based on or make particularly important use of data or results from that specific mission. Papers not included are likely to be those that make indirect use of the data from that particular mission or use the data only in an ancillary way. We have chosen to use queries that on the one hand ensure the inclusion in the samples of the majority of papers that make the greatest use of data from each mission, and on the other hand that are as simple as possible to ensure the greatest homogeneity when comparing samples from different missions.
The fraction of false positives is a strong function of the mission name. Searching for missions with familiar names such as “Fermi” or “Swift”, or “Dice” produces a large number of false positives. To limit them as much as possible, we added search keywords relevant to the missions. For example, in the search for “Fermi” we required the presence in the full text of the name of at least one of the Fermi instruments or targets: (facility:Fermi) OR (abs:Fermi AND (full:GRB OR full:”Fermi LAT” OR full:”FERMI/LAT” OR full:”Fermi GBM” OR full:”Fermi/GBM”)) year:2008-2024 property:refereed collection:astronomy. We stopped tailoring the query when the results converge to within a few %. We finally manually inspected the final paper list to look for obvious false positives.
@@
We use the ratio of the number of refereed papers to the number of years from operation, normalized for the total cost of missions in M\$ to parameterize mission productivity.
Following [26] we use the Research impact quotient $Riq=\sqrt{tori}/YSO$ to parameterize the impact of a mission. $YSO$ is again is again the number of years from operation. $tori$ is the total impact of the research given by $tori=\sum_{cit}1/ar$, where $cit$ is the collection of citations, $a$ is the number of authors of each citing paper and $r$ is the number of references of the citing paper. The $tori$ is defined as the amount of work devoted to exploitation of a given mission calculated in research papers. To assess the impact of the mission normalized for the mission cost, we use $Riq/\sqrt{M\$}$.
Acknowledgements This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. FF acknowledges useful discussions with Norbert Werner, Salvo Sciortino, Paolo Tozzi, Francesca Esposito.

### v2

@@ 5 Conclusions
(1) broaden the funding base. In addition to government money, scientists can fund their projects through philanthropic foundations, public-private partnerships, and private investors. Knowledge is a crucial space resource, and it is precisely the resource that basic science researchers are best at harvesting. Accurate Position, Navigation, and Timing, the largest business in today’s and also tomorrow’s space economy, moving hundreds billion dollars, is famously impossible without general relativity corrections. Ironically, Albert Einsten (the inventor of general relativity more than 100 years ago) and his heirs did not make a single dollar from this invention. More relevant to new space is the story of Riccardo Giacconi, 2002 Nobel Prize winner for the discovery of Cosmic X-ray sources and the Cosmic X-ray Background. When Giacconi joined Bruno Rossi of MIT in 1959 as a young researcher, he was hired by American Science and Engineering (AS&E), a company started the year before by Martin Annis and of which Bruno Rossi was chair of the board of directors. AS&E was working on NASA contracts, but was also very active in the market. They developed the first body scanner, later adopted in airports around the world. Money earned in the market allowed AS&E to let Rossi and Giacconi develop the strategies and instrumentation that led to the creation of X-ray astronomy and to a series of world-leading missions: Uhuru, Einstein Observatory, and Chandra [34]. This is one of the several possible scenarios that could link science and business. Today, the new space economy probably provides more opportunities than in the time of Rossi and Giacconi. Two broad areas seem particularly promising today and merit a mention: quantum technologies and resource prospecting on the Moon, Mars, and asteroids.
@@
(2) the ability of scientists to keep up with the innovations of the new space economy. Technology innovation proceeds through alternating phases of incremental and disruptive development. Business innovates through vertical integration and scale production, but also a service-oriented business model and a culture of innovation that manifest itself as openness to risk and iterative development. Giacconi again comes to our aid in highlighting these concepts. Soon after the discovery of a strong, seemingly uniform, X-ray background in 1962 Giacconi understood that to detect the sources responsible for the Cosmic X-ray background, an imaging quality of the order of an arcsecond was needed in X-rays, several thousands times better than then available. At those early times, it was difficult, if not impossible, to convince NASA to embark directly on a project to develop and launch an X-ray observatory with arcsecond capabilities. Instead, Giacconi imagined an iterative program aimed at consolidating the field and arriving at the arcsecond telescope (Chandra) through steps, Uhuru first, to provide the first all-sky survey of bright X-ray sources at the beginning of the 1970s, Einstein Observatory, to demonstrate in space the capabilities of X-ray optics, and finally the Advanced X-ray Astrophysics Facility, AXAF, renamed Chandra after launch in 1999. The new space economy offers the opportunity to perform iterative approach programs of this kind on much shorter timescales and at much lower cost.
@@
We have searched the Astrophysical Data System (ADS) and the SCOPUS database for refereed publications, citations, and the H index of a sample of recent space scientific missions funded by NASA and ESA. We included in the list a sample of cubesats realized by Universities and scientific institutes. The search was carried out using the simplest possible keywords, to be able to have a data sample as homogeneous as possible. As an example, the Gaia ADS Library produced by ESA includes 12.842 refereed papers since 2014. The library is obtained by merging the results of a complex and customized ADS query with a manually curated library. The authors claim a fraction of false positives $<$5%. The much simpler ADS query ((facility:Gaia OR abs:Gaia OR title:Gaia) year:2014-2024 property:refereed collection:astronomy) produces 6617 entries, of which $>$90% are present in the ESA Gaia ADS Library. A manual check on the papers missing in the ESA Gaia ADS Library shows that a large fraction are actually using Gaia data or results obtained from Gaia data. So we are confident that the fraction of false positives in our table is certainly smaller than 10% and probably smaller than 5%. Limiting the sample to papers that explicitly mention the name of the mission in the title, abstract, or facility keyword would include most papers that are based on or make particularly important use of data or results from that specific mission. Papers not included are likely to be those that make indirect use of the data from that particular mission or use the data only in an ancillary way. We have chosen to use queries that on the one hand ensure the inclusion in the samples of the majority of papers that make the greatest use of data from each mission, and on the other hand that are as simple as possible to ensure the greatest homogeneity when comparing samples from different missions.
The fraction of false positives is a strong function of the mission name. Searching for missions with familiar names such as “Fermi” or “Swift”, or “Dice” produces a large number of false positives. To limit them as much as possible, we added search keywords relevant to the missions. For example, in the search for “Fermi” we required the presence in the full text of the name of at least one of the Fermi instruments or targets: (facility:Fermi) OR (abs:Fermi AND (full:GRB OR full:”Fermi LAT” OR full:”FERMI/LAT” OR full:”Fermi GBM” OR full:”Fermi/GBM”)) year:2008-2024 property:refereed collection:astronomy. We stopped tailoring the query when the results converge to within a few %. Finally, we manually inspected the final paper list to look for obvious false positives.
@@
We use the ratio of the number of refereed articles to the number of years of operation, normalized for the total cost of missions in M\$ to parameterize mission productivity.
Following [35] we use the Research impact quotient $Riq=\sqrt{tori}/YSO$ to parameterize the impact of a mission. $YSO$ is again is again the number of years from operation. $tori$ is the total impact of the research given by $tori=\sum_{cit}1/ar$, where $cit$ is the collection of citations, $a$ is the number of authors in each article cited and $r$ is the number of references in the article cited. The $tori$ is defined as the amount of work devoted to exploitation of a given mission calculated in research papers. To assess the impact of the mission normalized for the mission cost, we use $Riq/\sqrt{M\$}$.
Acknowledgments This work was performed in part at the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452. FF acknowledges support from the INAF grant PRORIS. FF acknowledges useful discussions with Norbert Werner, Salvo Sciortino, Paolo Tozzi, Francesca Esposito.